# CheckEmbed: Effective Verification of LLM Solutions to Open-Ended Tasks

### v1

@@ 1 Introduction
Recent verification methods and their building blocks, such as SelfCheckGPT [33] and BERTScore [58], focus on individual fact checking and token- as well as sentence-level analysis. For example, SelfCheckGPT, albeit introduced recently, is an established mechanism for LLM verification. As one of its methods, it uses BERTScore to compare LLM answers. BERTScore is based on comparing token- or sentence-level embeddings. One obtains the embedding of each token within a sentence, and then uses a selected summary score to get the final sentence embedding. Such embeddings can be used to compare and reason about different tokens or sentences.
@@ 3 Scalability Analysis
In CheckEmbed, there are $k$ embeddings in total to construct, and thus $O(k^{2})$ similarity operations to be conducted. Next, one can apply BERTScore straightforwardly to two passages treated as long sentences. This means $O((st)^{2})=O(s^{2}t^{2})$ embedding comparisons have to be performed, resulting in a total of $O(k^{2}s^{2}t^{2})$ embedding comparisons and a total of $2k^{2}$ embedding constructions. Finally, SelfCheckGPT compares the sentences from a given selected reference sample to all samples collected. For each comparison, it uses BERTScore with $s^{2}$ sentences as candidates and $s^{2}$ sentences as a reference. This gives (using the above BERTScore formula) $O(ks^{2})$ embedding constructions and $O(ks^{2}s^{2}t)=O(ks^{4}t^{2})$ embedding comparisons.
@@ 4.3 Detecting Fine-Grained Hallucinations
While CheckEmbed is primarily targeted at verification of open-ended tasks, we also investigate whether CheckEmbed can be used to detect small fine-grained hallucinations, such as mistakes in individual facts. The results are in Figure 7 and the used prompts can be found in the Appendix A.1. The task analyzed is summarizing scientific articles. For each article considered, we generate a summary with no errors (labeled as “ground truth”), and we also ask the LLM to summarize these documents, while forcing deliberate small fact-level mistakes, from 0 to 10 mistakes per summary. We also compare to BERTScore and SelfCheckGPT. CheckEmbed is able to recognize when samples contains no errors, as illustrated by very large scores for GT. Moreover, interestingly, it can also recognize hallucinations after introducing a single error, as visible by no overlap between the GT and the consecutive data points. Finally, we can observe that the amount of low-confidence scores is somewhat increasing with the growing number of introduced errors. However, this increase only starts to be distinctive beyond 5 errors. The trends for BERTScore and SelfCheckGPT are similar, which illustrates that these baselines perform well for their intended use case.
Figure 7: Analysis of fine-grained hallucination verification of LLM answers (GPT-4o), details explained in Section 4.3.

### v2

@@ 1 Introduction
Recent verification methods and their building blocks, such as SelfCheckGPT [33] and BERTScore [58], focus on individual fact checking and token- as well as sentence-level analysis. For example, SelfCheckGPT, albeit introduced recently, is an established mechanism for LLM verification. As one of its methods, it uses BERTScore to compare LLM answers. In BERTScore, one first obtains contextual embeddings of all words in two compared sentences, computes cosine similarity scores between all pairs of embeddings from two sentences, and then greedily matches each token from one sentence to the most similar token in the other sentence. Weighted averages of these matchings are used to obtain precision, recall, and the final F1 score that enables comparing different sentences.
@@ 3 Scalability Analysis
In CheckEmbed, there are $k$ embeddings in total to construct, and $O(k^{2})$ similarity operations to be conducted.
Next, one can apply BERTScore straightforwardly to two passages treated as long sentences, each such passage consists of $st$ tokens. This means $O((st)^{2})=O(s^{2}t^{2})$ embedding comparisons have to be performed for any two passages (for each pair of compared sentences, one compares every pair of individual tokens/words), resulting in a total of $O(k^{2}s^{2}t^{2})$ embedding comparisons as this is done for $O(k^{2})$ pairs of LLM answers, and a total of $O(k^{2})$ embedding constructions.
Finally, SelfCheckGPT assesses a given LLM reply by comparing it to all sample replies collected. To simplify the following derivations, assume that in an individual comparison of two LLM replies, these replies consist of $s_{1}$ and $s_{2}$ sentences, respectively. Now, for each such comparison, SelfCheckGPT uses BERTScore, where the two input passages $x$ and $y$ to BERTScore consist of $s_{1}s_{2}$ sentences each, i.e., both passage $x$ and passage $y$ contain all the sentences from its corresponding LLM reply, repeated as many times as the number of sentences in the other LLM reply (this is condicted to enable comparing all sentences from each reply pairwise). This gives (using the above BERTScore formulae) $O(ks^{2})$ embedding constructions (there are $k$ LLM replies) and $O(ks^{2}s^{2}t^{2})=O(ks^{4}t^{2})$ embedding comparisons.
@@ 4.3 Detecting Fine-Grained Hallucinations
While CheckEmbed is primarily targeted at verification of open-ended tasks, we also investigate whether CheckEmbed can be used to detect small fine-grained hallucinations, such as mistakes in individual facts. The results are in Figure 7 and 8 and the used prompts can be found in the Appendix A.1. The task analyzed is summarizing scientific and legal articles. For each article considered, we generate a summary with no errors (labeled as “ground truth”), and we also ask the LLM to summarize these documents, while forcing deliberate small fact-level mistakes, from 0 to 10 mistakes per summary. CheckEmbed is able to recognize when samples contains no errors, as illustrated by very large scores for GT. Moreover, interestingly, it can also recognize hallucinations after introducing a single error, as visible by no overlap between the GT and the consecutive data points. Finally, we can observe that the amount of low-confidence scores is somewhat increasing with the growing number of introduced errors. However, this increase only starts to be distinctive beyond 5 errors. The trends for BERTScore and SelfCheckGPT are similar, which illustrates that these baselines perform well for their intended use case.
Figure 7: Analysis of fine-grained hallucination verification of LLM answers (GPT-4o) when summarizing scientific documents, details explained in Section 4.3.
Figure 8: Analysis of fine-grained hallucination verification of LLM answers (GPT-4o) when summarizing legal documents, details explained in Section 4.3.